We are given a wheel of readius "r" and mass "m" rotating about its axis with angular speed "v/r". If its radius is doubled, wat will be the change in net force acting on the particle at the rim ??


1) mass is constant.


2) density is constant.

Grade:12

1 Answers

vikas askiitian expert
13 years ago

since no external torque acting on the system so we can apply conservation of angular momentam...

Iw=I1w1 .........1

a)when mass is constant

eq 1 becomes

 mr2 w =m4r2w1

   w1=w/4 ..............1

the only force acting on the particle which is placed on rim is centrifugal force..

Fc=mw2 r               (initially)

Fc1=mw12 2r          (finally)

Fc1=mw2r/8=Fc/8

so force on particle becomes 1/8 times its initial value..

now solve other part in same manner